Question 12 Current research indicates that the distribution of the life expectancies of a certain protozoan is normal with a mean of 49 days and a standard deviation of 10.1 days. Find the probability that a simple random sample of 25 protozoa will have a mean life expectancy of 54 or more days. 0.2067 b) O0.0067 0.6897 0.9933 e) 0.0133
Answer #1

12)

Solution :

Given that ,

mean = \mu = 49

standard deviation = \sigma = 10.1

n = 25

\mu\bar x = 49 and

\sigma\bar x = \sigma / \sqrt n = 10.1 / \sqrt 25 = 2.02

P(\bar x? \geq 54) = 1 - P(\bar x\leq 54) = P((\bar x - \mu \bar x ) / \sigma \bar x\leq (54 - 49) / 2.02) = 1 - P(z \leq 2.473)

    Using standard normal table,

P(\bar x\geq 54) = 1 - 0.9933 = 0.0067

Probability = 0.0067

Option b) is correct .